Gina is 9 years younger than her sister Brigget. Two years ago Brigget was twice as old as Gina was. how old is Gina now?
Thank you for helping me

Answer by stanbon(48568)

Gina is 9 years younger than her sister Brigget. Two years ago Brigget was twice as old as Gina was. how old is Gina now?
NOW Data:
Brigget age = "x"
Gina age is "x-9"
TWO-YEARS AGO Data:
Brigget age = "x-2"
Gina age = "x-9-2"="x-11"
EQUATION:
Two years ago Brigget was twice as old as Gina.
x-2=2(x-11)
x-2=2x-22
x=20 (Brigget's age NOW.
x-9= 11(Gina's age NOW.
Cheers,
Stan H.
B